OpenStudy (anonymous):

find the area of a triangular window given the base and height .. b=2in h=3in please and thank you:)

OpenStudy (anonymous):

area = 1/2 * base * height = 1/2 * 2 * 3 = 3 sq ins

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

yw